12 concrete cylindrical pillars of a building are to be covered by aluminum
cladding. Each pillar has a diameter of 70 cm and height of 7m.
1. How much concrete is needed to construct the concrete pillars?
1
Expert's answer
2020-10-25T18:49:42-0400

"\text{The required volume }V\text{ of concrete is}"

"V=n*V_c\text{ where}"

"n=12\text{ number of columns } V_c \text{ volume of one column}"

"\text{according to the right cylinder volume formula }"

"V_c=\pi*r^2*h;\ h=7m;\ r=\frac{0.7}{2}=0.35m"

"V_c=\pi*0.35^2*7\approx2.69m^3"

"V= 12*2.69=32.28m^3"

Answer: "V=32.28m^3" required amount of concrete